OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (-2, 6) and has a slope of 4. 1.) y + 2 = 4(x - 6) 2.) y - 2 = 4(x + 6) 3.) y - 6 = 4(x + 2) 4.)y + 6 = 4(x - 2)

OpenStudy (anonymous):

y-6= 4(x-(-2)) y-6=4(x+2)

OpenStudy (anonymous):

thanks man :D